# Multimodal Clinical Trial Outcome Prediction with Large Language Models

### v3

@@ 2.4. Representation Augmentation and Consistency Loss
where $v_{i,k}^{0}\equiv[cls]_{k}$.
2.5. Integrating Multimodal Information with Mixture-of-Experts
@@
(9) $W_{i,k}=$ $\mathrm{Softmax}(\mathcal{C}(\oplus_{j\in\mathcal{J}}U_{i,j})*\gamma_{k})$
@@
(10) $\mathcal{L}_{c}=\frac{1}{N}$ $\sum_{i=1}^{N}\ell(\hat{y}_{i},y_{i}),$
$\mathrm{where}$ $\hat{y}_{i}=\mathcal{H}(U_{i})$
@@
(12) $\mathcal{L}=\mathcal{L}_{c}+\eta_{1}\mathcal{L}_{con}+\eta_{2}\mathcal{L}_{aux},$
where $\eta_{1}$ and $\eta_{2}$ are coefficients to balance these loss terms. The whole algorithm is illustrated in Alg. 1.
1
@@
2
3step 1. Transforming Multimodal Data into Natural Language Descriptions;
4 for $i\leftarrow 1$ to $N$ do
5       for $k\leftarrow 1$ to $K$ do
6             $x_{i,k}\leftarrow\mathrm{LLM}(p,\mathrm{Linearize}(x_{i,k}),s)$
7      $x_{i,0}\leftarrow\mathrm{LLM}(p,\mathrm{Linearize}(x_{i}),s)$
8step 2. Train LIFTED;
9 foreach minibatch $\mathcal{B}$ in dataset $\mathcal{D}$ do
10       for $i\leftarrow 1$ to batch size $b$ do
11             for $k\leftarrow 0$ to $K$ do
12                   $u_{i,k}^{t}\leftarrow\mathcal{E}_{k}(\mathcal{T}(z_{i,k}))$ $v_{i,k}^{t}\leftarrow\hat{u}_{i,k}^{t}$ as Equation 7 $U_{i,k}\leftarrow\mathcal{F}_{k}(\{u_{i,k}^{t}+pos^{t}\}_{t=0}^{T})$, $\tilde{U}_{i,k}\leftarrow\mathcal{G}(U_{i,k})\cdot\mathcal{R}(U_{i,k})$, $V_{i,k}\leftarrow\mathcal{F}_{k}(\{v_{i,k}^{t}+pos^{t}\}_{t=0}^{T})$, $\tilde{V}_{i,k}\leftarrow\mathcal{G}(V_{i,k})\cdot\mathcal{R}(V_{i,k})$
13            Fuse representations with MoE method as Equation (9)
14      Compute the losses $\mathcal{L}_{con}$, $\mathcal{L}_{c}$, $\mathcal{L}_{aux}$ and $\mathcal{L}$ following Equation (12)
15       Optimize the parameters $\theta$ of LIFTED by minimizing $\mathcal{L}$
@@
3. Experiments
In this section, we evaluate the performance of LIFTED  aiming to answer the following questions:
  • Q1: Compared to the existing methods with modal-specific encoders, can LIFTED achieve better performance with the unified transformer encoders?
  • Q2: Do the key components, including the multimodal data integration component and the representation augmentation component, of LIFTED boost the performance?
  • Q3: Does the Sparse Mixture-of-Experts framework route the similar information patterns in different modalities to the same expert models correctly?
  • Q4: Does the Mixture-of-Experts approach precisely measure the importance of different modalities for each patient?
3.1. Dataset Descriptions.
We evaluate our method and other baselines on the HINT dataset (Fu et al., 2022; Chen et al., 2024), which includes the information on diseases, the name, description, and SMILES string of drugs, eligibility criteria for each clinical trial record, the phase, and also, the trial outcome labels as success or failure covering Phases I, II and III trials. In our implementation, we incorporate all modalities, including disease, the name, description and SMILES string of drugs and criteria, totaling five modalities. Additionally, we include phase information when generating the natural language summarization for samples. The transform-based encoder for the SMILES string modality is pre-trained and the corresponding tokenizer is specifically designed for SMILES string data. However, all the other modalities are tokenized by a unified tokenizer and none of the other encoders are pre-trained. The HINT dataset contains 17,538 clinical trial records, with 1,787 trials in Phase I, 6,102 trials in Phase II, and 4,576 trials in Phase III (Fu et al., 2021). The detailed data statistics of the HINT dataset are shown in Table 1.
Table 1. The statistics of the HINT Datasets (Fu et al., 2021). # is short for the number of. The number of Trials is shown by the split of train/validation/test sets.
# Trials # Drugs # Diseases # patients/trial # Success # Failure
Phase I 1,044/116/627 2,020 1,392 45 1,006 781
Phase II 4,004/445/1,653 5,610 2,824 183 3,039 3,063
Phase III 3,092/344/1,140 4,727 1,619 1418 3,104 1,472
@@ 3. Experiments
Table 2. The clinical trial outcome performance (%) of LIFTED and baselines. The mean and standard deviations are calculated from 30 independent runs with different random seeds. : The results of the HINT and SPOT methods were obtained by running their released codes. The best results and second best results are bold and underlined, respectively. We observe that LIFTED consistently outperforms all other methods over all three phases.

### v4

@@ 2.3 Consistent Representation Augmentation
where we define $v_{i,k}^{0}\equiv[cls]_{k}$.
2.4 Integrating Multimodal Information
@@
$W_{i,k}=$ $\mathrm{Softmax}(\mathcal{C}(\oplus_{j\in\mathcal{J}}U_{i,j})*\gamma_{k})$ (9)
@@
$\mathcal{L}_{c}=\frac{1}{N}\sum_{i=1}^{N}\ell(\hat{y}_{i},y_{i}),\ \ \mathrm{where}\ \ \hat{y}_{i}=\mathcal{H}(U_{i})$ (10)
@@
$\mathcal{L}=\mathcal{L}_{c}+\eta_{1}\mathcal{L}_{con}+\eta_{2}\mathcal{L}_{aux},$ (12)
where $\eta_{1}$ and $\eta_{2}$ are hyperparameters to balance these loss terms. The whole algorithm is illustrated in Alg. 1.
@@
step 1. Transforming Multimodal Data into Natural Language Descriptions;
for $i\leftarrow 1$ to $N$ do
       for $k\leftarrow 1$ to $K$ do
             $x_{i,k}\leftarrow\mathrm{LLM}(p,\mathrm{Linearize}(x_{i,k}),s)$
      $x_{i,0}\leftarrow\mathrm{LLM}(p,\mathrm{Linearize}(x_{i}),s)$
step 2. Train LIFTED;
foreach minibatch $\mathcal{B}$ in dataset $\mathcal{D}$ do
       for $i\leftarrow 1$ to batch size $b$ do
             for $k\leftarrow 0$ to $K$ do
                   $u_{i,k}^{t}\leftarrow\mathcal{E}_{k}(\mathcal{T}(z_{i,k}))$
                   $v_{i,k}^{t}\leftarrow\hat{u}_{i,k}^{t}$ as Equation 7
                   $U_{i,k}\leftarrow\mathcal{F}_{k}(\{u_{i,k}^{t}+pos^{t}\}_{t=0}^{T})$,
                   $\tilde{U}_{i,k}\leftarrow\mathcal{G}(U_{i,k})\cdot\mathcal{R}(U_{i,k})$,
                   $V_{i,k}\leftarrow\mathcal{F}_{k}(\{v_{i,k}^{t}+pos^{t}\}_{t=0}^{T})$,
                   $\tilde{V}_{i,k}\leftarrow\mathcal{G}(V_{i,k})\cdot\mathcal{R}(V_{i,k})$
            Fuse representations with MoE method as (9)
      Compute the losses $\mathcal{L}_{con}$, $\mathcal{L}_{c}$, $\mathcal{L}_{aux}$ and $\mathcal{L}$
       Optimize the parameters $\theta$ of LIFTED
@@
3 Experiments
In this section, we evaluate the performance of LIFTED  aiming to answer the following questions: Q1: Compared to the existing methods with modal-specific encoders, can LIFTED achieve better performance with the unified transformer encoders? Q2: Do the key components, including the multimodal data integration component and the representation augmentation component, of LIFTED boost the performance? Q3: Does the Sparse Mixture-of-Experts framework route similar information patterns in different modalities to the same expert models correctly? Q4: Does the Mixture-of-Experts approach precisely measure the importance of different modalities for each patient?
@@ 3 Experiments
Table 1: The clinical trial outcome performance (%) of LIFTED and baselines on HINT dataset and CTOD dataset. The results are averaged on 30 independent runs with different random seeds. : The results of the HINT and SPOT methods were obtained by running their released codes. The best results and second best results are bold and underlined, respectively. We observe that LIFTED consistently outperforms all other methods over all three phases.
HINT CTOD
Phase I Trials Phase II Trials Phase III Trials Phase I Trials Phase II Trials Phase III Trials
Method PR F1 ROC PR F1 ROC PR F1 ROC PR F1 ROC PR F1 ROC PR F1 ROC
LR $50.0$ $60.4$ $52.0$ $56.5$ $55.5$ $58.7$ $68.7$ $69.8$ $65.0$ $55.6$ $58.5$ $51.5$ $56.0$ $63.6$ $56.7$ $75.0$ $73.8$ $51.9$
RF $51.8$ $62.1$ $52.5$ $57.8$ $56.3$ $58.8$ $69.2$ $68.6$ $66.3$ $58.3$ $70.0$ $54.0$ $60.8$ $70.9$ $56.0$ $75.1$ $85.1$ $50.7$
XGBoost $51.3$ $62.1$ $51.8$ $58.6$ $57.0$ $60.0$ $69.7$ $69.6$ $66.7$ $57.4$ $68.0$ $50.5$ $61.2$ $70.6$ $61.3$ $78.7$ $84.9$ $57.1$
AdaBoost $51.9$ $62.2$ $52.6$ $58.6$ $58.3$ $60.3$ $70.1$ $69.5$ $67.0$ $58.8$ $66.7$ $53.5$ $60.2$ $68.9$ $55.2$ $79.5$ $84.6$ $56.9$
kNN+RF $53.1$ $62.5$ $53.8$ $59.4$ $59.0$ $59.7$ $70.7$ $69.8$ $67.8$ $58.4$ $70.3$ $54.7$ $61.2$ $70.9$ $56.6$ $76.8$ $85.2$ $52.5$
FFNN $54.7$ $63.4$ $55.0$ $60.4$ $59.9$ $61.1$ $74.7$ $74.8$ $68.1$ $55.2$ $58.4$ $48.7$ $57.2$ $66.0$ $53.0$ $76.1$ $79.6$ $52.6$
MMF (early) $60.6$ $59.4$ $54.4$ $60.2$ $62.6$ $60.7$ $85.5$ $81.5$ $70.6$ $61.2$ $64.2$ $56.2$ $64.0$ $70.1$ $59.0$ $83.3$ $84.4$ $64.6$
MMF (late) $63.4$ $67.5$ $59.0$ $62.9$ $63.0$ $62.6$ $86.9$ $83.1$ $71.8$ $63.0$ $70.5$ $57.4$ $65.7$ $71.4$ $60.3$ $83.5$ $85.2$ $65.8$
HINT $58.4$ $68.2$ $62.1$ $59.1$ $63.9$ $62.8$ $85.9$ $80.9$ $70.8$ $63.4$ $71.0$ $57.6$ $64.6$ $71.2$ $60.6$ $81.8$ $85.7$ $60.8$
SPOT $69.8$ $68.4$ $64.6$ $62.6$ $64.3$ $63.0$ $81.7$ $81.0$ $71.0$ $66.8$ $70.0$ $62.3$ $64.5$ $71.8$ $58.8$ $83.2$ $77.6$ $67.5$
LIFTED 70.7 71.6 64.9 69.8 66.2 65.1 88.3 83.8 73.5 69.7 71.8 63.4 67.7 72.0 63.0 86.7 85.9 69.5
3.1 Experimental Setup
Dataset Descriptions. We evaluate our method and other baselines on the HINT dataset Fu et al. (2022); Chen et al. (2024) and CTOD dataset Gao et al. (2024), covering Phases I, II and III trials. More details of the HINT dataset and the CTOD dataset are shown in Appendix C.
Baselines.
We compare LIFTED with both machine learning methods and deep learning models, such as Feedforward Neural Network (FFNN) Shen et al. (2023), Multi Modal Fusion (MMF), HINT Fu et al. (2022); Wang et al. (2024), SPOT Wang et al. (2023b). More details of those baselines are presented in Appendix B.
Evaluation Metrics.
Following Fu et al. Fu et al. (2022) and Chen et al. Chen et al. (2024), we use F1 score, PR-AUC, and ROC-AUC to measure the performance of all methods. For all these metrics, higher scores indicate better performance.
Table 2: The clinical trial outcome prediction performance (%) of LIFTED and variants without certain key component. The best results are bold. LIFTED outperforms all variants, showcasing the effectiveness of our proposed components.
HINT
Phase I Trials Phase II Trials Phase III Trials
Method PR F1 ROC PR F1 ROC PR F1 ROC
LIFTED-aug $68.4$ $69.8$ $64.8$ $69.5$ $66.0$ $64.3$ $86.9$ $82.4$ $72.1$
LIFTED-aux $69.0$ $71.2$ $63.7$ $69.6$ $64.5$ $64.6$ $87.4$ $82.8$ $71.1$
LIFTED-LLM $68.5$ $70.8$ $64.0$ $69.7$ $64.9$ $65.0$ $86.7$ $82.7$ $70.8$
LIFTED-gating 69.9 71.3 64.9 69.7 65.5 65.0 87.0 82.7 72.4
LIFTED 70.7 71.6 64.9 69.8 66.2 65.1 88.3 83.8 73.5
CTOD
Phase I Trials Phase II Trials Phase III Trials
Method PR F1 ROC PR F1 ROC PR F1 ROC
LIFTED-aug $65.1$ $70.6$ $61.9$ $67.1$ $70.5$ $62.2$ $85.2$ $82.4$ $67.8$
LIFTED-aux $64.4$ $71.1$ $60.0$ $60.0$ $71.0$ $54.9$ $83.8$ $85.5$ $64.6$
LIFTED-LLM $65.1$ $71.4$ $58.5$ $66.3$ $71.1$ $61.2$ $83.9$ $85.4$ $65.3$
LIFTED-gating $67.1$ $71.4$ $60.8$ $65.1$ $70.4$ $61.9$ $85.0$ $85.8$ $68.8$
LIFTED 67.7 71.6 62.3 67.7 72.0 63.0 86.7 85.9 69.5
Table 3: Performance analysis of multimodal data integration. The best results and second best results are bold and underlined, respectively.
HINT
Phase I Trials Phase II Trials Phase III Trials
PR F1 ROC PR F1 ROC PR F1 ROC
Summarization $63.2$ $69.9$ $57.8$ $66.1$ $61.2$ $61.2$ $85.1$ $80.7$ $66.6$
Drugs $62.1$ $67.2$ $57.9$ $60.5$ $62.3$ $55.8$ $83.8$ $81.7$ $63.8$
Disease $65.3$ $67.3$ $59.7$ $68.0$ $59.9$ $62.4$ $86.0$ $80.5$ $69.1$
Description $55.5$ $71.3$ $50.2$ $55.5$ $0.0$ $50.0$ $74.9$ 85.7 $49.7$
SMILES $62.8$ $69.6$ $58.5$ $59.3$ $58.3$ $54.9$ $76.1$ $83.6$ $51.1$
Criteria $68.0$ $70.5$ $63.1$ $67.6$ $64.4$ $63.0$ $83.7$ $82.7$ $65.0$
All (LIFTED) 70.7 71.6 64.9 69.8 66.2 65.1 88.3 $83.8$ 73.5
CTOD
Phase I Trials Phase II Trials Phase III Trials
PR F1 ROC PR F1 ROC PR F1 ROC
Summarization $61.6$ $70.9$ $56.9$ $65.4$ $71.5$ $61.1$ $84.1$ $84.8$ $64.9$
Drugs $60.8$ $70.6$ $57.6$ $58.0$ $71.3$ $53.8$ $77.3$ $85.2$ $54.5$
Disease $65.2$ $70.2$ $61.3$ $68.3$ $71.1$ $62.6$ $85.3$ $85.5$ $68.4$
Description $56.6$ $70.7$ $52.3$ $56.0$ $71.3$ $52.4$ $76.5$ $85.7$ $52.9$
SMILES $61.6$ $70.6$ $56.0$ $58.0$ $71.5$ $53.1$ $75.1$ $85.6$ $50.8$
Criteria $60.5$ $71.2$ $55.8$ $59.9$ $71.2$ $53.9$ $75.7$ $84.6$ $52.5$
All (LIFTED) 67.7 71.6 62.3 67.7 72.0 63.0 86.7 85.9 69.5
3.2 Overall Performance
We conduct experiments to evaluate the performance of LIFTED on all three phases trails, compared to our baselines. The trial outcome prediction results of all models are reported in Table 1. We first observed that the deep learning-based methods and the methods designed for clinical trial outcome prediction outperform the machine learning based methods with a significant performance gap, especially on the HINT dataset, showcasing the powerful ability to extract critical information from different modalities in various formats of the deep learning encoders and those encoders specifically designed to extract representation hidden in the clinical trial records. This observation is not surprising, since the critical information of different modalities is represented in different ways, which is hard to extract for those traditional machine learning methods or those deep learning encoders that are not designed for clinical trial outcome prediction. Nevertheless, LIFTED consistently outperforms all other methods over all three phases, verifying its effectiveness in unifying different modalities and dynamically integrating them within the MoE.
3.3 Ablation Study
In this section, we perform comprehensive ablation studies to demonstrate the effectiveness of our key components, including the representation augmentation, the auxiliary loss and the modalities used to generate weights in the Mixture-of-Experts (MoE) framework.
  • LIFTED-aug: In LIFTED-aug, the representation augmentation component and the consistency loss are removed. Representations from different modalities are directly fed into the multimodal data integration component without the constraint of robustness to the noise in data.
  • LIFTED-aux: In LIFTED-aux, we remove the auxiliary loss component. Representations from different modalities are no longer required to make consistent predictions with the final representation.
3.4 Analysis of Multimodal Data Integration
We further analyze how multimodal data integration contributes to clinical outcome prediction. Here, we compare the performance of models using data from only one modality with LIFTED that integrates all those modalities. We report the results in Table 3. The results indicate that LIFTED outperforms almost all unimodal models, demonstrating the effectiveness of multimodal integration. In addition, the results also demonstrate that the drug description and the criteria modalities are the least and the most important modality, respectively, which is expected since the quality of recruited patients plays a crucial role in trial success Jin et al. (2017); Zhang et al. (2021).
3.5 Analysis of Sparse Mixture-of-Experts
Figure 3: The SMoE experts’ importance weights of our model predicting the knee osteoarthritis patient. Experts 6 and 7 play a crucial role in extracting common information patterns across modalities, while other experts specialize in a single specific modality.
In addition, we delve into an analysis of the Sparse MoE model to understand the performance enhancements obtained by it. Here, we select a knee osteoarthritis patient case. For each modality, the SMoE framework selects top-$3$ experts from a pool of $16$ experts with the highest weights. The weights of these selected SMoE experts are visualized in Figure 3. As expected, certain experts, such as 6 and 7, are consistently chosen across multiple modalities, indicating their pivotal role in extracting similar information patterns among different modalities. Furthermore, other experts demonstrate a more focused expertise, concentrating on one or two modalities. This demonstrates the effectiveness of the SMoE framework in both extracting similar information patterns across different modalities and capturing specialized information patterns within a single modality.
3.6 Case Study
Figure 4: The modality importance weights of our model predicting the type 2 diabetes mellitus patient. LIFTED pay more attention to the disease modality as expected, since type 2 diabetes mellitus is hard to cure.
In addition, we conduct a case study to analyze the contribution of each modality in clinical trial outcome prediction. Specifically, we analyze the result of a type 2 diabetes mellitus patient, who was inadequately controlled with metformin at the maximal effective and tolerated dose of metformin for at least 12 weeks. Since type 2 diabetes mellitus is hard to cure Chang et al. (2019), the model should pay attention to the name of the disease and predict the trial as failed, which is consistent with the behavior of our model. The modality importance weights are shown in Figure 4. As we expected, the attention weights of the disease modality are much higher than other modalities, which demonstrates that our model pays attention to the disease modality and predicts the trial correctly.
4 Related Works
Clinical Trial Outcome Prediction. Machine learning methods have been proven efficient on diverse tabular data prediction tasks, especially the clinical trial outcome prediction task, resulting in profound performances Chen et al. (2023); Yan et al. (2023). Recently, Fu et al. Fu et al. (2023) proposed a hierarchical interaction network employing different encoders to fuse multiple modal data and capture their correlations for trial outcome predictions; Wang et al. Wang et al. (2023b) clustered multi-sourced trial data into different topics, organizing trial embeddings for prediction. Wang et al. Wang et al. (2023a) converted clinical trial data into a format compatible description for prediction. However, converting all modalities into a single description poses significant challenges. This approach makes it difficult for the model to distinguish the unique information of each modality and necessitates external data to aid in differentiating these modalities. In contrast, LIFTED extracts representations for each modality separately and dynamically integrates them, providing a more effective way to preserve distinct characteristics of each modality.
Mixture-of-Experts. Mixture-of-Experts (MoE) is a special type of neural network whose parameters are partitioned into a series of sub-modules, called experts, functioning in a conditional computation fashion Jacobs et al. (1991); Jordan and Jacobs (1993). Recently, Shazzer et al. Shazeer et al. (2017) simplified the MoE layer by selecting a sparse combination of the experts, instead of all experts, to process input data, significantly reducing the computational cost and improving the training stability. To encourage specialization and decrease redundancy among experts Chen et al. (2022), Dai et al. Dai et al. (2022) pre-defined the expert assignment for different input categories, and Hazimeh et al. Hazimeh et al. (2021) advocated multiple, diverse router policies, facilitating the intriguing goals of SMoE is to divide and conquer the learning task by solving each piece of the task with adaptively selected experts Aoki et al. (2022); Mittal et al. (2022). To identify similar information patterns between different modalities and extract them with the same expert model, LIFTED follows the design of Sparse Mixture-of-Experts Shazeer et al. (2017), routing inputs to a subset of experts, dynamically selecting the experts instead of using a pre-defined assignment.
5 Conclusion
We introduce LIFTED, an approach that unifies multimodal data using natural language descriptions and integrates this information within a Mixture-of-Experts (MoE) framework for clinical trial outcome prediction. We employ noise-resilient encoders to extract representations from each modality, utilize a Sparse MoE framework to further dig the similar information patterns in different modalities, and introduce an auxiliary loss to improve the quality of modal-specific representations. Empirically, our LIFTED method demonstrates superior performance compared to existing approaches across all three phases of clinical trials, underscoring the effectiveness and potent representational capacity of natural language and highlighting the potential for a unified text modality to supplant diverse information modalities.
Appendix A Prompt
The whole prompt, including the system message, is demonstrated in Table A.1, and some examples are demonstrated in Table A.2.
Table A.1: Prompting.
System Message
You are a helpful assistant.
Prompting
Response
\$summarization_of_the_sample
Table A.2: Examples of Prompting.
Linearization Summarization
[…]
Appendix B Baselines
Many methods have been selected as baselines in our experiments, including both statistical machine learning and deep learning models. We use the same setups in Fu et al. Fu et al. [2022] and Wang et al. Wang et al. [2023b] for most of them.
  • Logistic regression (LR) Lo et al. [2019]; Siah et al. [2021]: logistic regression with the default hyperparameters implemented by scikit-learn Pedregosa et al. [2011].
  • Random Forest (RF) Lo et al. [2019]; Siah et al. [2021]: similar to logistic regression, the random forest is also implemented by scikit-learn with the default hyperparameters Pedregosa et al. [2011].
  • XGBoost Rajpurkar et al. [2020]; Siah et al. [2021]: An implementation of gradient-boosted decision trees optimized for speed and performance.
  • Adaptive boosting (AdaBoost) Fan et al. [2020]: an adaptive boosting-based decision tree method implemented by scikit-learn Pedregosa et al. [2011].
  • k Nearest Neighbor (kNN) + RF Lo et al. [2019]: a combined model using kNN to imputate missing data and predicting by random forests.
  • Feedforward Neural Network (FFNN) Tranchevent et al. [2019]: a feedforward neural network that uses the same feature as HINT Fu et al. [2023]. The FFNN contains three fully-connected layers with hidden dimensions of 500 and 100, as well as a rectified linear unit (ReLU) activation layer to provide nonlinearity.
  • Multi-Modal Fusion (MMF): This technique amalgamates multi-modal data to arrive at a final prediction, employing both early fusion and late fusion strategies. In the early fusion approach, various modal inputs are first concatenated before being fed into the prediction model. Conversely, in the late fusion variant, multiple prediction models are employed on each modal input, and the ultimate prediction is derived through fusion techniques, such as voting, which integrates predictions from each modality.
  • HINT Fu et al. [2022]: several key components are integrated with HINT, including a drug molecule encoder utilizing MPNN algorithm, a disease ontology encoder based on GRAM, a trial eligibility criteria encoder leveraging BERT, and also, a drug molecule pharmacokinetic encoder, surplus a graph neural network to capture feature interactions. After the interacted features are encoded, they are fed into a prediction model for accurate outcome predictions.
  • SPOT Wang et al. [2023b]: SPOT contains several steps. Firstly, trial topics are identified to group the diverse trial data from multiple sources into relevant trial topics. Subsequently, trial embeddings are produced and organized according to topic and timestamp to construct organized clinical trial sequences. Finally, each trial sequence is treated as a separate task, and a meta-learning approach is employed to adapt to new tasks with minimal modifications.
Appendix C Dataset Descriptions
Table C.3: The statistics of the HINT Dataset and the CTOD Dataset. The number of Trials is shown by the split of train/validation/test sets.
HINT
# Trials # Drugs # Diseases # Success # Failure
Phase I 1,044/116/627 2,020 1,392 1,006 781
Phase II 4,004/445/1,653 5,610 2,824 3,039 3,063
Phase III 3,092/344/1,140 4,727 1,619 3,104 1,472
CTOD
# Trials # Drugs # Diseases # Success # Failure
Phase I 1,012/149/627 1,638 913 1,129 659
Phase II 3,950/501/1,653 5,003 2,254 3,949 2,156
Phase III 3,075/363/1,140 3,863 1,533 3,643 941
The HINT dataset Fu et al. [2022]; Chen et al. [2024] and CTOD dataset Gao et al. [2024] include the information on diseases, the name, description, and SMILES string of drugs, eligibility criteria for each clinical trial record, the phase, and also, the trial outcome labels as success or failure covering Phases I, II and III trials. The HINT dataset contains 17,538 clinical trial records, with 1,787 trials in Phase I, 6,102 trials in Phase II, and 4,576 trials in Phase III Fu et al. [2021]. We utilize parts of the CTOD dataset to test LIFTED’s performance, with 1,788 trials in Phase I, 6,105 trials in Phase II, and 4,584 trials in Phase III, for a total of 12,477 trials. Gao et al. [2024]. The detailed data statistics of the HINT dataset and the CTOD dataset are shown in Table C.3.
In our implementation, we incorporate all modalities, including disease, the name, description and SMILES string of drugs and criteria, totaling five modalities. Additionally, we include phase information when generating the natural language summarization for samples. The transform-based encoder for the SMILES string modality is pre-trained and the corresponding tokenizer is specifically designed for SMILES string data. However, all the other modalities are tokenized by a unified tokenizer and none of the other encoders are pre-trained.
Appendix D Hyperparameter Settings
We follow the settings of most hyperparameters in HINT Fu et al. [2022]. The models are trained for a total of 5 epochs using a mini-batch size of 32 on one NVIDIA 4090 GPUs, which will take up to 2 hours. We employ the AdamW optimizer Loshchilov and Hutter [2017] with a learning rate of $3\times 10^{-4}$, $\beta$ values of $(0.9,0.99)$, and a weight decay of $1\times 10^{-2}$ with a CosineAnnealing learning rate scheduler.
Appendix E Detailed Results
Table E.4: The clinical trial outcome performance (%) of LIFTED and baselines on HINT dataset and CTOD dataset. The mean and standard deviations are calculated from 30 independent runs with different random seeds. : The results of the HINT and SPOT methods were obtained by running their released codes. The best results and second best results are bold and underlined, respectively. We observe that LIFTED consistently outperforms all other methods over all three phases.
HINT